a) Using Cayley’s theorem, find the permutation group to which a cyclic group of

order 12 is isomorphic. (4)

b) Let τ be a fixed odd permutation in . S10 Show that every odd permutation in S is 10

a product of τ and some permutation in . A10 (2)

c) List two distinct cosets of < r > in , D10 where r is a reflection in . D10 (2)

d) Give the smallest n ∈ N for which An is non-abelian. Justify your answer.


1
Expert's answer
2021-05-20T11:39:12-0400

a) Given a cyclic group G = <a> of order 12 generated by the element 'a’ of G, we may take the 'permutation group’ as the cyclic subgroup H of S"_{12}" generated by the cycle (1,2,3,4,5,6,7,8,9,10,11,12) of length 12. This (or a conjugate of it) is what we obtain as per the proof of the Cayley's Theorem.

We may also find isomorphic copies of G as subgroups of groups S"_n" for all n "\ge"7. For example the product (1,2,3,4)•(5,6,7) of disjoint cycles in S"_7" has order LCM(4,3)=12 and hence generates a cyclic subgroup of S"_7" , isomorphic to G.


d)

"A_n" has order "n!\/2"  and is not abelian if "n\ge4"

Answer: "n=4"


c)

Let "D_{10}=\{x^iy^i|x^{10}=y^2=1,xy=yx^9\}"

two cosets: "\{x,x^6\},\ \{x^2,x^7\}"


b)

"A_{10}" the subset of even permutations in "S_{10}".

"\tau" is product of odd number of transpositions in "S_{10}" ; permutation in "A_{10}" is product of even number of transpositions in "S_{10}" . So, result of product of "\tau" and permutation in "A_{10}" is:

(odd number of transpositions in "S_{10}" )+(even number of transpositions in "S_{10}" )=

=(odd number of transpositions in "S_{10}" ), that is odd permutation in "S_{10}".